OpenStudy (firejay5):

Simplify. 5 times the square root of - 24 * 3 times the square root of -18

OpenStudy (anonymous):

Oh wait

OpenStudy (firejay5):

It's 5 times the square root of -24 X 3 times the square root of -18

OpenStudy (anonymous):

Yea I read that wrong. Our equation is 3*(sqrt-18) first. Sorry

OpenStudy (firejay5):

Our first problem is 5 times the square root of -24

OpenStudy (anonymous):

No actually if the problem is: 5 times the square root of - 24 * 3 times the square root of -18. It is 5*(sqrt-24)(3*sqrt18)

OpenStudy (anonymous):

Since this is our equation I chose to start with 3*sqrt of 18

OpenStudy (anonymous):

Do you understand?

OpenStudy (firejay5):

\[5\sqrt{-24} * 3\sqrt{-18}\]

OpenStudy (anonymous):

Okay. Well in either case need to know how to square negatives. To do this we need to use the imaginary unit i. The equation for that will be i=(-1)^(1/2). Now we substitute i with our negative: -24=-24*-1 Understand so far?

OpenStudy (firejay5):

Okay, what I have so far is \[5\sqrt{4} * \sqrt{6} * \sqrt{-1} = 5\sqrt{-24} (and) 3\sqrt{9} * \sqrt{2} * \sqrt{-1} = 3 \sqrt{-18}\]

OpenStudy (firejay5):

I know the answer is -180 times the square root of 3. I just need to know how to work it out

OpenStudy (anonymous):

Hm.. Well I used the imaginary unit and simplified 5sqrt-24 to 10i sqrt6. Then for 3sqrt -18 I got 9i sqrt2. Then I multiplied to get -180 sqrt 3. So it'd be 5*sqrt 24 simplified by 3*sqrt18 simplified. (Sorry for the wait)

OpenStudy (anonymous):

\[5\sqrt{-24} * 3\sqrt{-18}\]\[5i\sqrt{4*6} * 3i\sqrt{6*3}\]\[15i^2\sqrt{4}\sqrt{6}\sqrt{6}\sqrt{3}\]\[-15*2*6\sqrt{3}\]\[-180\sqrt3\]

OpenStudy (firejay5):

what if I\[\sqrt{9} \sqrt{2}\] used

OpenStudy (firejay5):

@ChmE

OpenStudy (anonymous):

Then the sqrt of 6 wouldn't be able to multiply together

OpenStudy (firejay5):

what would I get then

OpenStudy (anonymous):

\[15i^2\sqrt{4}\sqrt{6}\sqrt{9}\sqrt{2}\]\[15i^2*2*3\sqrt{6}\sqrt{2}\]\[90i^2\sqrt{12}\]\[90i^2\sqrt{4}\sqrt3\]\[-180\sqrt3\]Looks like the same answer just more work

OpenStudy (firejay5):

it's okay as long as it's the same answer